ΔH° is -58.03 kJ and ΔS° is -176.6J/K. What is the value of ΔG° at 298K? Is the reaction spontaneous?
1
Expert's answer
2020-05-01T14:22:47-0400

Solution.

∆G0=∆H0-T∆S0.

∆S0= -176,6 J/K= -0,1766 kJ/K.

∆G0298= -58,03-298*(-0,1766) =-5,4032 kJ.

Since ∆G0298< 0 the reaction is spontaneous.

Answer: ∆G0298 = -5,4032 kJ. The reaction is spontaneous.